# Uncovering thermodynamic origin of counterflow and coflow instabilities in miscible binary superfluids

### v1

@@ IV Holographic binary superfluids
We then move to the gravity description of a two-component superfluid in two spatial dimensions. More details are given in appendix A. The system involves two complex scalar operators $\mathcal{\hat{O}}_{I}$ which carry the same charge $e$ under a global $U(1)$ symmetry. [^6] Thus there will be only one conserved density, which is a special limit of the hydrodynamic and GPE analyses of the previous sections. Below a critical temperature $T_{c}$, the scalar operators develop a nonzero expectation value spontaneously breaking the $U(1)$ symmetry and driving the system into a superfluid phase. The holographic system in (3+1)-dimensional asymptotically anti-de Sitter (AdS) spacetime involves a bulk dynamical $U(1)$ gauge field $A_{\mu}$ and the scalar operators $\mathcal{\hat{O}}_{i}$ are mapped to two bulk scalar fields $(\Psi_{1},\Psi_{2})$ carrying charge $e$ under $A_{\mu}$. For definiteness we will choose two identical scalar fields with the potential $V(\Psi_{1},\Psi_{2})=m^{2}(|\Psi_{1}|^{2}+|\Psi_{2}|^{2})+\frac{\nu}{2}|\Psi_{1}|^{2}|\Psi_{2}|^{2}$ with a mass for $(\Psi_{1},\Psi_{2})$ correspond to the dual scalar operators having the scaling dimension $\Delta=2$. The inter-component coupling $\nu$ characterizes the interaction between two components of superfluids and determines the miscibility of the superfluids An et al. (2024a). For the homogeneous case with $\nu>0$, only one of the superfluid condensations can develop spontaneously below the critical temperature $T_{c}$ (immiscible binary superfluids), while $\nu<0$ gives miscible binary superfluids, which is the focus of this work.
@@
At finite temperature, each of the two species contains both superfluid and normal components. This appears to be a rather complicated problem. To avoid these complications, without loss of generality we consider the case where the fluctuations of the temperature and the normal fluid velocity are frozen, and meanwhile, the dynamics of momentum and energy are decoupled from the charge sector. This corresponds to the probe limit in the gravity description for which the back-reaction of matter content to the geometry is neglected. More precisely, the background is given by the Schwarzschild AdS black brane (see Appendix A for details on the bulk action and the metric we use). After solving the bulk EoMs, we can read off all relevant observables by the standard holographic dictionary, including the temperature $T$, the superfluid velocity $\bm{v}_{I}^{s}\equiv(v_{Ix}^{s},v_{Iy}^{s})$ for each superfluid condensation $\mathcal{O}_{I}\equiv\langle\mathcal{\hat{O}}_{I}\rangle$. The gravity dual thus provides a first-principle description of superfluid dynamics.
@@
The phase-separated binary superfluids have been studied recently in An et al. (2024a, b); An and Li (2024) in holography, where rich dynamics were uncovered. Here we are interested in the holographic miscible binary superfluids, which provides a good platform for checking the hydrodynamic framework in Section II. The construction of the stationary configuration for miscible binary superfluids is provided in Appendix B.
@@ IV.1 Dynamical instability from linear analysis
Without loss of generality, we assume that the superfluids move along the $y$ direction with the velocity $v_{y}$. More precisely, the counterflow case corresponds to $\bm{v}_{1}^{s}=-\bm{v}_{2}^{s}=(0,v_{y})$, and the coflow one has $\bm{v}_{1}^{s}=\bm{v}_{2}^{s}=(0,v_{y})$. Since the instability will typically develop fastest along the flow of the superfluids, we consider the wave vectors parallel to the superfluid velocity, i.e. $\bm{k\cdot x}=k\,y$. The full set of linearized EoMs can be found in Appendix D, which results in a generalized eigenvalue problem that can be solved numerically. In Figure 1, we give a representative example for the counterflow instability (upper panel) and the coflow instability (lower panel), where we show the dominant QNMs with respect to the wave number $k$. While the spectral patterns are quite different, one finds that a dynamical instability will develop beyond a critical velocity for both counterflow and coflow cases.
@@
The spectrum of QNMs for the case with $\nu=0$ and $v_{y}=0$ was studied in Amado et al. (2013) where the scalar fields transform as a doublet, i.e. a global $SU(2)$ symmetry. It was found that there exist two kinds of gapless modes. One is the sound mode of the superfluid, and the dispersion relations is linear in the hydrodynamic limit $\omega_{\mathrm{I}}=\pm(v_{s}k+\bar{b}k^{2})-i\Gamma k^{2}$ which is classified as the type $\mathrm{I}$ Goldstone boson Hidaka (2013) and is due to the spontaneous breaking of $U(1)$ symmetry. The other presents quadratic dispersion relation at small $k$, $\omega_{\mathrm{II}}=\pm bk^{2}-ick^{2}$, and is classified as the type $\mathrm{II}$ Goldstone boson. This mode appears because of spontaneous breaking of the bulk global $SU(2)$ symmetry. When $\nu\neq 0$, the $SU(2)$ symmetry is explicitly broken. But there is a residual $U(1)\times U(1)$ symmetry. Therefore, one branch of type $\mathrm{II}$ Goldstone boson becomes gapped, while the other branch remains gapless but becomes a dissipative, quadratic mode. This is clearly shown in Figure 2, from which there are two modes with linear dispersion relation (red line), one mode with quadratic dispersion relation (blue line) and a gapped mode (black line).
@@
We find that this pattern is true for all parameters of our holographic binary superfluids, as long as $\nu<0$. And this is also consistent with the result of (14) from hydrodynamics of binary miscible superfluids (which does not incorporate the dynamics of the gapped mode).
Figure 2: The QNM spectrum for holographic miscible binary superfluids with $T/T_{c}=0.677$, $\nu=-0.1$ and $v_{y}=0$. Left panel shows real part of QNMs while right panel shows imaginary part. Red lines are sound modes. Blue line and black line correspond to type $\mathrm{II}$ Goldstone boson when $\nu=0$. At nonzero $\nu$, we expect one branch of this mode to otain an imaginary gap, which is indeed what we observe.
@@
IV.1.1 Dispersion relation for counterflow case
@@
In the counterflow case for which $\bm{v}_{1}^{s}=-\bm{v}_{2}^{s}=(0,v_{y})$, the instability first presents itself in one of the sound modes of (25). More precisely, above the critical velocity $v_{c1}$, the sound speed becomes purely imaginary ($v_{s}^{2}<0$) and one of them obtains a positive imaginary part, see the red lines in the upper panel of Figure 3. There is another critical velocity $v_{c2}$, above which the unstable mode becomes gapped at $k=0$, as shown by the black line in the lower panel of Figure 3.
@@
Figure 3: Spectrum of QNMs at low momentum for counterflow case. Top: Spectrum of QNMs for superfluid velocity $v_{y1}=0.283$. Bottom: Spectrum of QNMs for superfluid velocity $v_{y2}=0.314$. Red lines are sound modes, blue line is diffusive mode and black line is gapped mode. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$. Note $v_{c1}<v_{y1}<v_{c2}$ and $v_{y2}>v_{c2}$. The unstable mode for $v_{y1}$ is sound mode, while for $v_{y2}$, the unstable mode is the gapped mode. For reference, the two critical velocities at $T/T_{c}=0.677$ and $\nu=-0.2$ are numerically found to be $v_{c1}=0.251$, $v_{c2}=0.308$.
@@
Figure 4: Top: Sound velocity of sound mode $v_{s}$ for different superfluid velocities $v_{y}$ in counterflow case. Bottom: Susceptibility $\chi_{++}$ and free energy density difference $\Delta f=f_{\mathrm{binary}}-f_{\mathrm{single}}$. The vertical dashed lines denote critical velocities $v_{c1}$ and $v_{c2}$. Above $v_{c1}$, $\chi_{++}$ changes sign, the sound velocity becomes purely imaginary, and one sound mode becomes unstable. Above $v_{c2}$, $\Delta f$ becomes positive. When $\Delta f>0$, the binary superfluid phase is globally thermodynamically unstable, and the gapped mode becomes unstable with a positive imaginary part at $k=0$. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$.
@@
As predicted by the hydrodynamics of binary superfluids in Section II, the onset of thermodynamic instability for is given by (17). Since $\chi_{+-}$=0 for both counterflow and coflow case, one then has the criterion $\chi_{++}=0$ or $\chi_{--}=0$. We find that the former one $\chi_{++}=\delta^{2}f/\delta^{2}v^{s}_{+}|_{v^{s}_{-}}=0$ explains precisely the dynamical instability we observe in the counterflow case, while $\chi_{--}$ does not change sign. According to the holographic dictionary, the free energy density $f$ is computed from the on-shell bulk action with Euclidean signature times temperature. Then we can extract $\chi_{++}$ by calculating the free energy density for different $v^{s}_{+}$ while keeping $v^{s}_{-}$ fixed. See Appendix C.1 for more details. In Figure 4, we plot the sound velocity $v_{s}$ and $\chi_{++}$ for different superfluid velocities $v_{y}$.  [^8] In Figure 4, the sound velocity is extracted from QNMs directly by $v_{s}=\lim_{k\rightarrow 0}\omega/k$. One can also extract sound velocity from susceptibilities according to (15). We have verified that results from these two methods agree within numerical error, and the discrepancy is given in Appendix E. The two critical velocities $v_{c1}$ and $v_{c2}$ are denoted by vertical dashed lines. We see the the first critical velocity corresponds exactly to $\chi_{++}=0$. Note that $\chi_{n+}\sim v^{s}_{+}$. Therefore $\chi_{n+}=0$ for the counterflow case, where $v^{s}_{+}=0$. Then from (15), the sound velocity is $v_{s}=\sqrt{\chi_{++}/\chi_{nn}}$. One can see the sound velocity should become pure imaginary when $\chi_{++}$ changes sign, which is exactly what happens to the sound velocity above the first critical velocity (see the dashed blue line of Figure 4).
@@
We point out that the fastest growing mode for above instability develops at a finite wave number. Interestingly, as $v_{y}$ is increased, $\chi_{++}$ becomes positive again, for which hydrodynamics of Section II suggests no instability. Surprisingly, we find that, beyond the second critical velocity $v_{c2}$, the system does become unstable at $k=0$ while retaining a positive $\chi_{++}$. This is a global instability associated with the gapped mode becoming unstable, due to order competition between the two components.  [^9] Early studies on competing holographic orders can be found, e.g. in Basu et al. (2010); Cai et al. (2013). Note that there exist two kinds of broken phases: the single superfluid phase and the binary superfluid phase. When $v_{y}>v_{c2}$, the binary superfluid phase becomes less thermodynamically favored and order competing between the two components triggers a global instability. We can verify this by comparing the free energy of the two phases. By calculating the difference of free energy density between the two superfluid phases $\Delta f=f_{\mathrm{binary}}-f_{\mathrm{single}}$, we find that $\Delta f$ becomes positive exactly when $v_{y}>v_{c2}$, see the black line in the bottom panel of Figure 4. Therefore, there will be a first order phase transition from the binary superfluid phase to the single superfluid phase. Moreover, as shown in Figure 4, $\chi_{++}$ diverges at $v_{c2}$. This instability is indeed thermodynamic origin but beyond the scope of our hydrodynamic description.
IV.1.2 Dispersion relation for coflow case
@@
We now turn to the coflow case with $\bm{v}_{1}^{s}=\bm{v}_{2}^{s}=(0,v_{y})$. We first point that the boost symmetry of the GPEs (18) removes the coflow instability for the binary superfluids without relative velocity. Indeed, at zero temperature where the GPE applies, there is no normal component and any nonzero superfluid velocity can be boosted back using a Galilean transformation to a state where the superfluid is at rest. The holographic setup enjoys relativistic symmetry and is invariant under Lorentz boosts. Nevertheless, at finite temperature, states at nonzero superfluid velocity cannot be boosted back to a state where all components are at rest, since there is a preferred rest frame for the normal component, defined by setting the normal velocity to be $u^{\mu}=(1,0,0)$. Indeed, the interface instability for coflow immiscible binary superfluids that is absent in GPE has been recently observed in holographic superfluids An et al. (2024b).
@@
In sharp contrast to the case from GPE (18), we do observe coflow instavility from the holographic miscible binary superfluids. More precisely, we find two critical velocities denoted as $v^{\prime}_{c1}$ and $v^{\prime}_{c2}$. Comparing with the counterflow case, we find that the binary superfluid phase is always thermodynamically preferred for the coflow case, and thus there is no global instability.
@@
Figure 5: Spectrum of QNMs for coflow case. Top: Spectrum of QNMs for superfluid velocity $v^{\prime}_{y1}=1.257$. Bottom: Spectrum of QNMs for superfluid velocity $v^{\prime}_{y2}=2.199$. Red lines are sound modes, blue line is the diffusive mode and black line is the gapped mode. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$. Note $v^{\prime}_{c1}<v^{\prime}_{y1}<v^{\prime}_{c2}$ and $v^{\prime}_{y2}>v^{\prime}_{c2}$. For $v^{\prime}_{y1}$ only the diffusive mode is unstable, while for $v^{\prime}_{y2}$, one of the sound mode also becomes unstable. For reference, the two critical velocities at $T/T_{c}=0.677$ and $\nu=-0.2$ are numerically found to be $v^{\prime}_{c1}=0.691$, $v^{\prime}_{c2}=2.136$.
@@
As depicted in the upper panel of Figure 5, the instability first occurs from the diffusive mode of (25) when the superfluid velocity is above $v^{\prime}_{c1}$. Meanwhile, the sound and gapped modes are stable. As the superfluid velocity increases further to above $v^{\prime}_{c2}$, the other instability shows up from one of the sound modes, see red lines in the lower panel of Figure 5. We find that these two critical velocities exactly correspond to $\chi_{--}=0$ and $\chi_{++}=0$, respectively. For the coflow case case with $v^{s}_{-}=0$, the attenuation is $\Gamma_{0}=-\chi_{--}\delta_{1}/4$ from (15). Therefore, when $\chi_{--}$ changes sign, the diffusive mode becomes unstable. Moreover, when $\chi_{++}$ changes sign, the sound velocity becomes negative, and meanwhile $\Gamma$ becomes positive. To see this, we expand the sound velocity and attenuation in (15) around $\chi_{++}=0$. Without loss of generality, we assume $\chi_{n+}>0$. Then one of the sound velocity and attenuation would be
$v_{s}=$ $\delta\chi_{++}/2\chi_{n+}\,,$ (26)
$\Gamma=$ $-\frac{\delta\chi_{++}}{16}(\delta_{1}+\delta_{2}(v_{+}^{s})^{2}+\frac{4\beta v_{+}^{s}}{\chi_{n+}}+\frac{4\sigma}{\chi_{n+}^{2}})\,.$
Negative $v_{s}$ would signal excitation of modes with negative energy, and thus implies energetic instability.
@@
Figure 6: Top: Sound velocity $v_{s}$ of the unstable sound mode and attenuation $\Gamma_{0}$ of the diffusive mode for different superfluid velocities $v_{y}$ in coflow case. Bottom: susceptibilities $\chi_{++}$ and $\chi_{--}$. The vertical dashed lines denote the critical velocities $v^{\prime}_{c1}$ and $v^{\prime}_{c2}$, given by $\chi_{--}=0$ and $\chi_{++}=0$ respectively. Above $v^{\prime}_{c1}$, the diffusivity changes sign and the diffusive mode becomes unstable. Above $v^{\prime}_{c2}$, the sound mode also becomes unstable, with its velocity and attenuation both changing sign. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$.
@@
As shown in Figure 6, the above features from hydrodynamics are exactly what happens at the two critical velocities (vertical lines) in the holographic binary superfluids. At $v_{c1}^{\prime}$, both $\chi_{--}$ and $\Gamma_{0}$ vanish and change sign. At the larger critical velocity $v_{c2}^{\prime}$, $\chi_{++}$, $v_{s}$ and $\Gamma$ change their signs. Note that $\chi_{--}$ is always negative when $v_{y}>v_{c1}^{\prime}$. As we mentioned in Appendix D, there are two sets of solutions for the coflow case $(u_{1},v_{1})=\pm(u_{2},v_{2})$. By checking the eigenfunctions numerically, we find that the sound modes correspond to $(u_{1},v_{1})=(u_{2},v_{2})$, while the quadratic mode and the gapped mode correspond to $(u_{1},v_{1})=-(u_{2},v_{2})$. Note $(u_{I},v_{I})|_{z=0}$ are just perturbation of condensations on the boundary. This makes the physical meaning of the two kinds of coflow instability clear. For the first instability, perturbations of the two condensates are in the opposite phase, i.e. the two components have relative motion. For the second instability, perturbations are in the same phase, i.e. the two components move as a whole. In this sense, we can consider the second instability corresponding to the generalization of Landau instability in binary superfluids, which is due to both superfluids moving relative to normal component.
IV.2 Nonlinear stage for counterflow and coflow instability

### v2

@@ IV Holographic binary superfluids
We now move to the gravity description of a two-component superfluid in two spatial dimensions. More details about the holographic model are given in appendix C. The field theory system involves two complex scalar operators $\mathcal{\hat{O}}_{I}$, which carry the same charge $e$ under a global $U(1)$ symmetry. [^12] Thus there will be only one conserved density, which is a special limit of the hydrodynamic and GPE analyses of the previous sections. Below a critical temperature $T_{c}$, the scalar operators develop a nonzero expectation value, which breaks the $U(1)$ symmetry spontaneously and drives the system into a superfluid phase. The dual (3+1)-dimensional gravity system (the bulk) is defined on an asymptotically AdS spacetime. It includes a bulk dynamical $U(1)$ gauge field $A_{\mu}$ and two bulk complex scalar fields $(\Psi_{1},\Psi_{2})$ carrying charge $e$ under $A_{\mu}$, dual to the field theory scalar operators $\mathcal{\hat{O}}_{I}$. For definiteness, we will choose two identical scalar fields with the potential $V(\Psi_{1},\Psi_{2})=m^{2}(|\Psi_{1}|^{2}+|\Psi_{2}|^{2})+\frac{\nu}{2}|\Psi_{1}|^{2}|\Psi_{2}|^{2}$. [^13] Though the scalar fields are only directly coupled through the density-density coupling term, they are also indirectly coupled through interactions with gauge field. Therefore, current-current entrainment is naturally incorporated in our model. To see this explicitly, one can refer to Figure 6 and notice that $2N|_{v_{+}=v_{-}=0}=(\chi_{++}-\chi_{--})|_{v_{+}=v_{-}=0}$. We pick the value $m^{2}L^{2}=-2$, so that the dual scalar operators have the scaling dimension $\Delta_{\Psi}=2$. We do not expect our results to differ qualitatively for other choices of the mass. There is an inter-component coupling $\nu$, which characterizes the interaction between the two superfluid components and determines whether the components are miscible An et al. (2024a). For the homogeneous case with $\nu>0$, only one of the superfluid components can condense below the critical temperature $T_{c}$ (immiscible binary superfluids), while $\nu<0$ gives miscible binary superfluids, which is the focus of this work.
@@
At finite temperature, in general each of the two species contains both superfluid and normal components. However, the essential physics of the counterflow and coflow instabilities we are after can be accessed by freezing out the normal velocity and temperature fluctuations. This implies that the dynamics of momentum and energy are decoupled from the charge and superfluid sectors in the field theory. This corresponds to the so-called probe limit in the gravity description, where the backreaction of the matter sector on the metric is neglected. More precisely, the background is given by the Schwarzschild AdS black brane (see Appendix C for details on the bulk action and the metric we use). After solving the bulk EoMs, we can read off all the relevant observables thanks to the standard holographic dictionary, such as the temperature $T$ or the superfluid velocity $\bm{v}_{I}^{s}\equiv(v_{Ix}^{s},v_{Iy}^{s})$ for each superfluid condensate $\mathcal{O}_{I}\equiv\langle\mathcal{\hat{O}}_{I}\rangle$. This gravity dual provides a first-principle description of superfluid dynamics.
@@
Phase-separated binary holographic superfluids have been studied recently in An et al. (2024a, b); An and Li (2024), and rich dynamics was found. In this work, we are interested in miscible binary holographic superfluids, which provide a good platform for checking the hydrodynamic framework developed in Section II.1 and the onset of instabilities. The construction of the stationary configuration for miscible binary superfluids is provided in Appendix D.
@@ IV.1 Dynamical instability from linear analysis
Without loss of generality, we assume that the superfluid velocity $v_{y}$ is aligned along the $y$ direction. More precisely, the counterflow case corresponds to $\bm{v}_{1}^{s}=-\bm{v}_{2}^{s}=(0,v_{y})$, and the coflow case to $\bm{v}_{1}^{s}=\bm{v}_{2}^{s}=(0,v_{y})$. Since the instability will typically develop fastest along the flow of the superfluids, we restrict our analysis to wavevectors parallel to the superfluid velocity, i.e. $\bm{k\cdot x}=k\,y$.
The full set of linearized EoMs can be found in Appendix F, which results in a generalized eigenvalue problem that can be solved numerically. In Figure 1, we give a representative example for the counterflow instability (upper panel) and the coflow instability (lower panel), where we show the dominant QNMs with respect to the wave number $k$. While the spectral patterns are quite different, one finds that a dynamical instability will develop beyond a critical velocity for both the counterflow and coflow cases.
@@
The spectrum of QNMs for the case with $\nu=0$ and $v_{y}=0$ was studied in Amado et al. (2013) where the scalar fields transform as a doublet, i.e. there is a global $SU(2)$ symmetry (systems with such a symmetry have been studied in the context of binary superfluids for some time Manakov (1974)). In Amado et al. (2013), the holographic analysis found that there exist two kinds of gapless modes. The first are the superfluid sound modes, and their dispersion relations is linear in the hydrodynamic regime $\omega_{\mathrm{I}}=\pm(v_{s}k+\bar{b}k^{2})-i\Gamma k^{2}$. These modes are classified as type $\mathrm{I}$ Goldstone boson Hidaka (2013) and due to the spontaneous breaking of the $U(1)$ symmetry. The second type of gapless modes presents a quadratic dispersion relation at small $k$, $\omega_{\mathrm{II}}=\pm bk^{2}-ick^{2}$, and are classified as type $\mathrm{II}$ Goldstone bosons. These modes appear because the bulk global $SU(2)$ symmetry is spontaneously broken. Instead, when $\nu\neq 0$, the $SU(2)$ symmetry is explicitly broken, leaving a residual $U(1)\times U(1)$ symmetry. Therefore, one branch of type $\mathrm{II}$ Goldstone boson becomes gapped, while the other branch remains gapless but becomes a purely dissipative, quadratic mode. This is clearly shown in Figure 2, which shows there are two modes with real linear dispersion relations at small $k$ (red lines), one purely imaginary gapless mode with quadratic dispersion relation (blue line) and a purely imaginary gapped mode (black line).
@@
We find that this pattern is true for all parameters of our holographic binary superfluids, as long as $\nu<0$. This is also consistent with the dispersion relations (19) derived from hydrodynamics of binary miscible superfluids (which does not incorporate the dynamics of the gapped mode).
Figure 2: The QNM spectrum for holographic miscible binary superfluids with $T/T_{c}=0.677$, $\nu=-0.1$ and $v_{y}=0$. The left panel shows real part of QNMs while the right panel shows imaginary part. The red lines are the superfluid sound modes. The blue and black lines correspond to the type $\mathrm{II}$ Goldstone boson when $\nu=0$. At nonzero $\nu$, we expect one branch of this mode to otain an imaginary gap, which is indeed what we observe.
@@
IV.1.1 Dispersion relation for the counterflow case
@@
In the counterflow case $\bm{v}_{1}^{s}=-\bm{v}_{2}^{s}=(0,v_{y})$, the instability appears first in one of the sound modes of (36). More precisely, above the critical velocity $v_{c1}$, the speed of sound of one of the modes becomes purely imaginary ($v_{s}^{2}<0$), and this mode obtains a positive imaginary part, see the red lines in the upper panel of Figure 3. There is another critical velocity $v_{c2}$, above which the unstable mode becomes gapped at $k=0$, as shown by the black line in the lower panel of Figure 3.
@@
Figure 3: QNM spectrum at low momentum for the counterflow case. Top: QNM spectrum for superfluid velocity $v_{y1}=0.283$, so that $v_{c1}<v_{y1}<v_{c2}$. Bottom: QNM spectrum for superfluid velocity $v_{c2}<v_{y2}=0.314$. The red lines are the sound modes, the blue line is the diffusive mode and the black line is the gapped mode. The relevant parameter values are $T/T_{c}=0.677$, $\nu=-0.2$. The unstable mode for $v_{y1}$ is the sound mode, while for $v_{y2}$, the unstable mode is the gapped mode. For reference, the two critical velocities at $T/T_{c}=0.677$ and $\nu=-0.2$ are numerically found to be $v_{c1}=0.251$, $v_{c2}=0.308$.
@@
Figure 4: Top:Velocity $v_{s}$ of the sound mode for different superfluid velocities $v_{y}$ in the counterflow case. Bottom: Susceptibility $\chi_{++}$ and free energy density difference $\Delta f=f_{\mathrm{binary}}-f_{\mathrm{single}}$. The vertical dashed lines denote the critical velocities $v_{c1}$ and $v_{c2}$. Above $v_{c1}$, $\chi_{++}$ changes sign, the sound velocity becomes purely imaginary, and one sound mode becomes unstable. Above $v_{c2}$, $\Delta f$ becomes positive. When $\Delta f>0$, the binary superfluid phase is globally thermodynamically unstable, and the gapped mode becomes unstable with a positive imaginary part at $k=0$. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$.
@@
As predicted by the hydrodynamics of binary superfluids in Section II.1, the onset of the dynamical instability is given by (22). Since $\chi_{+-}$=0 for both the counterflow and coflow cases, the criterion reduces to $\chi_{++}=0$ or $\chi_{--}=0$. We find that the former one $\chi_{++}=0$ explains precisely the dynamical instability we observe in the counterflow case, while $\chi_{--}$ does not change sign. According to the holographic dictionary, the free energy $F$ at fixed superfluid velocity is computed from the on-shell bulk action with Euclidean signature times temperature. Then we can extract $\chi_{++}$ by calculating the free energy density for different $v^{s}_{+}$ while keeping $v^{s}_{-}$ fixed. See Appendix E.1 for more details. In Figure 4, we plot the sound velocity $v_{s}$ of the unstable mode and $\chi_{++}$ for different superfluid velocities $v_{y}$. [^15] In Figure 4, the sound velocity is extracted from the QNMs directly by $v_{s}=\lim_{k\rightarrow 0}\omega/k$. One can also compute the sound velocity from the susceptibilities according to (20). We have verified that the results from these two methods agree within numerical error, and the discrepancy is given in Appendix G. The two critical velocities $v_{c1}$ and $v_{c2}$ are denoted by vertical dashed lines. The first critical velocity occurs exactly when $\chi_{++}=0$. Then from (23), we expect the sound velocity to become purely imaginary when $\chi_{++}$ changes sign, which is exactly what happens above the first critical velocity (see the dashed blue line of Figure 4).
@@
We point out that the fastest unstable mode develops at a finite wave number. Interestingly, as $v_{y}$ is increased, $\chi_{++}$ becomes positive again. In this case, the hydrodynamics of Section II.1 predicts no instability. Nevertheless, we find that, beyond the second critical velocity $v_{c2}$, the system does become unstable at $k=0$ despite a positive $\chi_{++}$. This is in fact a global instability associated with the gapped mode becoming unstable, due to order competition between the two components.  [^16] Early studies on competing holographic orders can be found e.g. in Basu et al. (2010); Cai et al. (2013). Note that there exist two kinds of broken phases: the single superfluid phase and the binary superfluid phase. When $v_{y}>v_{c2}$, the binary superfluid phase becomes less thermodynamically favored, and order competing between the two components triggers a global instability. We can verify this by comparing the free energy of the two phases. By calculating the free energy density difference between the two superfluid phases $\Delta f=f_{\mathrm{binary}}-f_{\mathrm{single}}$, we find that $\Delta f$ becomes positive exactly when $v_{y}>v_{c2}$, see the black line in the bottom panel of Figure 4. Therefore, there will be a first-order phase transition from the binary superfluid phase to the single superfluid phase. Moreover, as shown in Figure 4, $\chi_{++}$ diverges at $v_{c2}$. This instability is also of thermodynamic origin but beyond the scope of our hydrodynamic description, as it is a global rather than a local instability.
IV.1.2 Dispersion relations for the coflow case
@@
We now consider the coflow case, $\bm{v}_{1}^{s}=\bm{v}_{2}^{s}=(0,v_{y})$. We first recall that the boost symmetry of the GPEs (29) removes the coflow instability for binary superfluids without relative velocity. Indeed, at zero temperature where the GPE applies, there is no normal component and any nonzero superfluid velocity can be boosted back using a Galilean transformation to a state where the superfluid is at rest.
Instead, the holographic setup enjoys relativistic symmetry and is invariant under Lorentz boosts. However, at finite temperature, states with a nonzero superfluid velocity cannot be boosted back to a state where all components are at rest, since there is a preferred rest frame for the normal component, defined by setting the normal velocity to be $u^{\mu}=(1,0,0)$. This means that the interface coflow instability for immiscible binary superfluids, which is absent in the GPE, can be observed in holographic superfluids An et al. (2024b).
@@
In the coflow case, we find two critical velocities denoted as $v^{\prime}_{c1}$ and $v^{\prime}_{c2}$. Comparing with the counterflow case, we find that in the coflow case the binary superfluid phase always has the lowest free energy, and thus there is no global instability towards a single superfluid phase. Thus, we expect both critical velocities should follow from the thermodynamic criterion (22).
@@
Figure 5: QNM spectrum for the coflow case. Top: QNM spectrum for a superfluid velocity $v^{\prime}_{y1}=1.257$, such that $v^{\prime}_{c1}<v^{\prime}_{y1}<v^{\prime}_{c2}$. Bottom: QNM spectrum for a superfluid velocity $v^{\prime}_{y2}=2.199$, such that $v^{\prime}_{y2}>v^{\prime}_{c2}$. The red lines are the superfluid sound modes, the blue line is the diffusive mode and the black line is the gapped mode. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$. At $v^{\prime}_{y1}$, only the diffusive mode is unstable, while at $v^{\prime}_{y2}$, one of the sound mode also becomes unstable. For reference, the two critical velocities at $T/T_{c}=0.677$ and $\nu=-0.2$ are numerically found to be $v^{\prime}_{c1}=0.691$, $v^{\prime}_{c2}=2.136$.
@@
As depicted in the upper panel of Figure 5, the diffusive mode is the first to become unstable when the superfluid velocity crosses over $v^{\prime}_{c1}$. Meanwhile, the sound modes and the gapped mode remain stable. As the superfluid velocity increases further to above $v^{\prime}_{c2}$, the other instability shows up from one of the sound modes, see the red lines in the lower panel of Figure 5. We find that these two critical velocities exactly correspond to $\chi_{--}=0$ and $\chi_{++}=0$, respectively. For the coflow case case with $v^{s}_{-}=0$, the attenuation is $\Gamma_{0}=-\chi_{--}\delta_{1}/4$, (25). Therefore, when $\chi_{--}$ changes sign, the diffusive mode becomes unstable. Moreover, when $\chi_{++}$ changes sign, the velocity of one of the sound modes (27) becomes negative (28), and its attenuation becomes positive. A negative sound speed signals modes with negative energy are excited, and thus implies an energetic instability, which is also accompanied by a dynamical instability due to the positive imaginary part.
@@
Figure 6: Top: Sound velocity $v_{s}$ of the unstable sound mode and attenuation $\Gamma_{0}$ of the diffusive mode for different superfluid velocities $v_{y}$ in the coflow case. Bottom: susceptibilities $\chi_{++}$ and $\chi_{--}$. The vertical dashed lines denote the critical velocities $v^{\prime}_{c1}$ and $v^{\prime}_{c2}$, given by $\chi_{--}=0$ and $\chi_{++}=0$ respectively. Above $v^{\prime}_{c1}$, the diffusivity changes sign and the diffusive mode becomes unstable. Above $v^{\prime}_{c2}$, the sound mode also becomes unstable, with its velocity and attenuation both changing sign. Relevant parameters are $T/T_{c}=0.677$, $\nu=-0.2$.
@@
As shown in Figure 6, the values for the critical velocities found from the QNM analysis exactly match the thermodynamic prediction from hydrodynamics. At $v_{c1}^{\prime}$, both $\chi_{--}$ and $\Gamma_{0}$ vanish and change sign. At the larger critical velocity $v_{c2}^{\prime}$, $\chi_{++}$, $v_{s}$ and $\Gamma$ change sign. Note that $\chi_{--}$ is always negative when $v_{y}>v_{c1}^{\prime}$. As we explain in Appendix F, in the coflow case, there are two sets of solutions $(u_{1},v_{1})=\pm(u_{2},v_{2})$ to the eigenvalue problem governing the QNMs. By checking the eigenfunctions numerically, we find that the sound modes correspond to $(u_{1},v_{1})=(u_{2},v_{2})$, while the quadratic mode and the gapped mode correspond to $(u_{1},v_{1})=-(u_{2},v_{2})$. Sincec $(u_{I},v_{I})$ are the fluctuations of the condensates at the boundary, This makes the physical meaning of the two kinds of coflow instability clear. For the first instability, the two condensates are fluctuating with opposite phases, i.e. the two components have relative motion. For the second instability, the condensates are fluctuating in phase, i.e. the two components move as a whole. In this sense, the second instability can be interpreted as the generalization of the Landau instability to binary superfluids, which is due to both superfluids moving relative to the normal component. The Landau instability in single-species superfluids was studied in Amado et al. (2014); Lan et al. ; Goutéraux et al. (2023); Areán et al. (2024).
IV.2 Nonlinear time evolution of the counterflow and coflow instabilities